Calculate the mass of the precipitate formed by mixing 20 g of an 8% solution of ferrous sulphate with a sufficient quantity of sodium hydroxide solution.


1
Expert's answer
2021-03-22T06:37:59-0400

FeSO4 + 2NaOH "\to" Fe(OH)2 + Na2SO4

1 mol of FeSO4 give 1 mole of Fe(OH)2

20 g of 8 %of ferrous sulphate

8 g of ferrous sulphate in 100 g

4 g of ferrous sulphate in 20 g

moles of ferrous sulphate = 4/151.908 = 0.026 mol

Mass of precipitate Fe(OH)2 = 0.026×89.86

= 2.33g